I am too confused with the concept of relative velocity.I will be thankful if you could explain me the concept with the help of the below problem.
A bus is beginning to move with an acceleration of 1metre/second sqaure.A boy who is 48m behind the bus starts running with constant speed of 10m/s.The time when the boy can catch the bus?
A)8sec
B)10sec
C)12sec
D)14sec.

Grade:11

2 Answers

basit ali
12 years ago

ITS ANS. IS  B) 10sec.  when the bus has accn. then its vel.(u) is 0 . and at that time boy has vel.(v)= 10m/s. then apply;         v=u+at.;or 10=1*t;or t=10s.                      and here relative velocity is also 10. plz reply to my answer.

nikhil roy
12 years ago

THE ANSWER IS 8SEC.